OpenStudy (bekkah323):

Solve the system of equations by graphing. Then classify the system. x+y=6 x-y=4

OpenStudy (anonymous):

Use https://www.desmos.com/calculator

pooja195 (pooja195):

x+y=6 <---e1 x-y=4 <---e2 We need to isolate x so that we can sub in the value add y to both sides x-y=4+y x=y+4 Now sub in x into e1 now that we know the value of x y+4+y=6 solve for y